# arXiv:2603.00192v1[cs.LG] 27 Feb 2026

### v1

@@ 5 Discussion
Individual-level prediction instability is an often overlooked barrier to the deployment of machine learning in healthcare. It may not be immediately obvious why instability in individual-level predictions is problematic in practice. First, the variance of a risk estimate is often overlooked. Second, the only working optimization routine may also be contributing significantly to the instability. One might argue that variability across model fits is an inherent feature of predictive modeling, and that different models will naturally sort individuals differently. However, not only may the dependency of the predictions on the learning algorithm be a useful quality to know and communicate, but the sensitivity of individual predictions represents an additional and largely unexamined source of uncertainty that merits explicit consideration alongside standard measures of generalization performance. While statistical learning literature has long recognized that flexible models can be more sensitive to arbitrary changes in the training data (Efron, 2004) and hence can exhibit overly optimistic training performance, our findings highlight a distinct, complementary concern. This concern being that flexible models can also be more sensitive to arbitrary changes to the learning pipeline. We do not claim that optimization-induced instability is equivalent to overfitting in the traditional sense; rather, it represents a form of procedural arbitrariness where a patient’s treatment eligibility is determined as much by the random initialization of the model as by their clinical data.
From a methodological perspective, our results expose the fragility of evaluating predictive models based on a single fitted instance. Relying on a single training run obscures meaningful variability; likewise, relying on a single risk estimate mistakes a stochastic snapshot for a definitive clinical truth. This is particularly problematic for overparameterized models where the loss landscape contains many nearly-equivalent local minima that nonetheless produce divergent clinical recommendations. In extreme cases, individual predictions may be effectively determined by random initialization rather than stable structure in the data. For certain modeling architectures and optimization procedures, a reported risk score is often better understood as a realization from a wide distribution of possible scores.
The evaluation framework we propose provides a practical way to characterize the sensitivity of the model to the learning pipeline (e.g., data, hyperparameters, computational settings). This framework can be especially valuable when comparing model families that are otherwise indistinguishable under conventional performance metrics. Our results advocate for a stability-aware interpretation of Occam’s razor: when out-of-sample performance is comparable, the more constrained model, such as a logistic regression model, can produce more reliable predictions compared to a neural network. By preferring models with greater individual-level stability, we mitigate the risk of procedural arbitrariness and ensure that clinical decisions are rooted in patient data rather than algorithmic noise. In high-stakes healthcare, where a "decision flip" can alter a patient’s care trajectory, individual-level instability should be elevated from a secondary consideration to a primary criterion for model selection.
Finally, these findings help contextualize the "uselessness" described by Markowetz (2024) and the persistent skepticism among clinical stakeholders. Skepticism toward AI is often dismissed as a lack of technical literacy (Kumar et al., 2025) or resistance to change (Sobaih and others, 2025). Our results suggest a more grounded justification: even when an overparameterized model appears accurate in aggregate, its individual-level predictions may be functionally inconsistent. From a clinician’s perspective, a model that flips its recommendation for the same patient across retrainings is not a reliable tool. Caution toward deployment, therefore, is not a rejection of innovation, but a rational response to a form of uncertainty that has been largely invisible under conventional validation practices. By making this variability explicit, our diagnostics provide a necessary bridge between technical validation and the rigorous reliability required for clinical care.
Box 1: Practitioner Checklist for Clinical Prediction Stability To operationalize the assessment of individual-level reliability, we propose the following checklist for evaluating whether a model’s procedural consistency meets the standards required for high-stakes healthcare deployment: What is the dispersion of an individual’s risk score across repeated retrainings? How frequently do perturbations in the learning pipeline cause a patient’s classification to flip? Are there specific subpopulations or risk strata where prediction instability is disproportionately high?
@@ 6 Conclusion
This work demonstrates that clinical prediction models with indistinguishable aggregate performance can nonetheless exhibit substantial, hidden variability at the level of individual predictions. We show that this instability arises not only from variations in training data but also from stochastic elements inherent to the learning pipeline, such as random initialization and optimization noise. Our results indicate that for high-capacity models, these optimization-induced fluctuations can be as impactful as resampling the entire training dataset.
To make this phenomenon operational, we introduced two model-agnostic diagnostics: empirical prediction interval width (ePIW) and empirical decision flip rate (eDFR). Ultimately, we find that individual-level instability can be valuable to incorporate for overparameterized clinical model evaluations. By incorporating stability diagnostics into standard validation workflows, practitioners can move beyond aggregate metrics to ensure that model-assisted decisions are robust to the arbitrary elements of the machine learning pipeline. When multiple model specifications achieve comparable out-of-sample performance, our results suggest that a selection strategy centered on procedural consistency can be warranted: prioritizing more constrained models that yield stable individual-level predictions may be a necessary step toward building the trust and reliability required for high-stakes clinical deployment.

### v2

@@ 5 Discussion
Individual-level prediction instability is an often overlooked barrier to the deployment of machine learning in healthcare. It may not be immediately obvious why instability in individual-level predictions is problematic in practice. First, the variance of a risk estimate is often overlooked. Second, the only working optimization routine may also be contributing significantly to the instability. One might argue that variability across model fits is an inherent feature of predictive modeling, and that different models will naturally sort individuals differently. However, not only may the dependency of the predictions on the learning algorithm be a useful quality to know and communicate, but the sensitivity of individual predictions represents an additional and largely unexamined source of uncertainty that merits explicit consideration alongside standard measures of generalization performance. Statistical learning literature has long recognized that flexible models can be more sensitive to arbitrary changes in the training data (Efron, 2004), our findings highlight a distinct, complementary concern. This concern being that flexible models can also be more sensitive to arbitrary changes to the learning pipeline. We do not claim that optimization-induced instability is equivalent to overfitting in the traditional sense, rather, it represents a form of procedural arbitrariness where a patient’s treatment eligibility is determined as much by the random initialization of the model as by their clinical data.
Our results reveal a tradeoff that mirrors the classical bias-variance decomposition but extends it into the domain of procedural consistency. More constrained models, such as logistic regression, exhibit low instability but may carry greater systematic bias relative to the true risk function, particularly in settings where the data-generating process is nonlinear. More flexible models can achieve lower bias at the cost of substantially greater instability. In settings where simpler models achieve comparable aggregate performance, as in our experiments, the instability cost of additional flexibility may not be justifiable. When there is a meaningful performance gap favoring the more flexible model, practitioners must weigh the gains in accuracy against the losses in individual-level consistency, taking into account the clinical consequences of inconsistent recommendations at the specific decision threshold in use.
The evaluation framework we propose provides a practical means of characterizing model sensitivity to the learning pipeline across the full range of model complexity. Stability diagnostics should be incorporated into routine model validation workflows rather than treated as optional post-hoc analyses. To support this, we propose the following checklist for evaluating whether a model’s procedural consistency meets the standards required for high-stakes clinical deployment:
  • What is the dispersion of an individual’s predicted risk across repeated retrainings, and does it vary across the risk spectrum?
  • How frequently do perturbations in the learning pipeline cause a patient’s clinical classification to flip?
  • Are there specific subpopulations or risk strata where prediction instability is disproportionately high, and do these coincide with clinically actionable thresholds?
  • Which specific modeling choices (e.g., architecture, optimizer, initialization) systematically induce greater instability, and can more stable alternatives achieve comparable predictive performance?
Finally, these findings help contextualize some skepticism among clinical stakeholders toward AI-assisted decision making (Markowetz, 2024). Such skepticism is often attributed to a lack of technical literacy (Kumar et al., 2025) or resistance to change (Arvai et al., 2025). Our results suggest a more principled justification: individual-level predictions may be functionally inconsistent across retrainings depending on the model capacity. It is worth noting that the preferred balance between stability and flexibility may vary by clinical task. In settings where consistent, reliable predictions are highly valued (e.g., routine screening) more constrained models that sacrifice flexibility for procedural consistency may be preferable. In contrast, tasks that require discriminating among hard cases, where a more flexible model’s ability to capture complex structure may outweigh its instability, may warrant a different tradeoff. Our framework does not prescribe a single answer, but rather makes the stability dimension of this choice explicit so that practitioners can reason about it deliberately rather than inadvertently.
@@ 6 Conclusion
This work demonstrates that clinical prediction models with indistinguishable aggregate performance can nonetheless exhibit substantial, hidden variability at the level of individual predictions. This is variability that arises not only from the training data but from stochastic elements of the learning pipeline itself, such as random initialization and optimization noise. For high-capacity models, these optimization-induced fluctuations can be as impactful as resampling the entire training dataset, extending the classical understanding of model complexity and prediction stability into the overparameterized regime where the stochastic nature of optimization introduces an additional and largely unaudited source of uncertainty. To quantify this phenomenon, we introduced two model-agnostic diagnostics, empirical prediction interval width (ePIW) and empirical decision flip rate (eDFR). Taken together, these findings establish individual-level reliability as an essential and complementary criterion for clinical model evaluation.